### v1

@@ 4.3.1 RQ7: How do cloning practices vary across different types of asset files in VR software?
Insight 7.  In VR software, the cloning behavior of asset files varies with complexity. Complex files like scenes have lower clone levels, simpler ones like materials have higher, and prefab files fluctuate with development style.
@@ 5.1.1 For Researchers
We recommend that researchers should give thorough consideration to the distinctions between different software types and general software when conducting studies on code cloning, ensuring a more comprehensive exploration of these differences. In particular, VR software necessitates deeper exploration of the cloning phenomena in its specialized asset and configuration files. The current study utilizes large language model technology to examine cloning behaviors in these files, hoping to provide significant insights for future investigations in this domain.
5.1.2 For VR Software Developers
We recommend that VR software developers carefully address the issue of code cloning throughout the development process to mitigate potential software vulnerabilities and minimize maintenance challenges. This includes using design patterns and coding paradigms that are suitable for the project, adopting modular design, avoiding direct copying and pasting of code blocks or files, and utilizing static analysis tools for cloning evaluation during the software development process. Specifically, for the unique asset files in VR software, developers should also consider applying similar design patterns and coding paradigms, such as using variants or instantiations, to avoid direct duplication of asset files. Since these files cannot be detected by traditional detection tools, the large model detection method used in this study could be applied. By employing these strategies, developers can effectively prevent or mitigate code cloning issues, enhancing the software’s maintainability, scalability, and overall quality.
5.1.3 For Third-party Library Designers
VR software often includes cloned function classes within third-party libraries, which is particularly common in libraries with cross-platform capabilities. We recommend using specific techniques to mitigate these cloning phenomena. For example, platform abstraction layers or interface abstract classes can be used to encapsulate specific implementations, conditional compilation directives can compile different code blocks for different platforms, dependency injection, and factory classes can enable dynamic implementation, or cross-platform frameworks (e.g., OpenXR) can be employed. Additionally, projects should select third-party libraries that best meet their needs; for example, if the project only needs to run on a specific hardware platform, a lightweight third-party library suitable for that platform can be chosen.
5.2 Future Work
Future research on the topic of this study could be approached from several directions. First, regarding the asset file types in VR software, our research does not cover all file types. Some files, such as audio, images, textures, and others stored in specialized formats, are not thoroughly investigated. This could be a key direction for future studies. Secondly, our evaluation of source code and serialized asset file cloning in VR software is conducted separately. A possible future direction would be to integrate these two aspects into a unified system for a comprehensive assessment. Finally, although we use the best available cloning detection tools for source code, with the advancement of AI technologies, future research could explore the use of large language models for detecting both source code and serialized asset files together. A potential solution could involve designing a large language model agent specifically for cloning behaviors in VR software, which could also be integrated with the aforementioned comprehensive evaluation of both file types.

### v2

@@ 4.3.1 RQ7: How do cloning practices vary across different types of asset files in VR software?
Insight 8.In VR software, the cloning behavior of asset files varies with complexity. Complex files like scenes have lower clone levels, simpler ones like materials have higher, and prefab files fluctuate with development style.
@@ 5.1.1 For Researchers
This study yields several actionable implications for researchers studying software cloning in VR development. First, the strong correlation between project size and cloning suggests that scalable clone detection tools are necessary for large VR projects, particularly in the gaming and education domains where cloning is most prevalent. Second, the distinct reliance on serialized asset files rather than source code highlights the importance of developing specialized detection methods tailored to asset-based cloning. Third, security-focused research should particularly investigate inter-version cloning patterns that facilitate vulnerability propagation. Finally, to advance clone analysis in VR software, researchers should focus on creating context-aware, multi-faceted metrics that capture the unique structural and behavioral features of VR systems. This effort should be supported by a comprehensive evaluation framework combining empirical validation, contextual experimentation, and cross-scenario applicability testing to ensure both precision and practical impact.
5.1.2 For VR Software Developers and Practitioners
This study also yields several actionable implications for VR software developers and practitioners. First, employ robust design patterns and modular architectures to facilitate code reuse and minimize reliance on copy-paste practices, particularly in large-scale VR projects where code and asset cloning is more prevalent.
Second, the integration of clone detection tools into development workflows is essential for the early identification and refactoring of duplicated content, thereby mitigating long-term maintenance burdens in increasingly complex systems.
Third, asset management should be treated with the same rigor as source code maintenance. Rather than generating slight variations of materials or textures, developers are encouraged to establish a centralized asset repository and adopt reusable instances or prefabs to ensure consistency and reduce redundancy across scenes.
Fourth, development practices should align with the characteristics of the chosen programming language. In C#, for instance, developers can leverage ScriptableObjects and inheritance to encapsulate reusable logic, whereas in C or C++, repetitive tasks should be delegated to optimized libraries or engine-level modules instead of being implemented repeatedly.
Fifth, evaluate third-party libraries carefully and consider alternatives with lower cloning levels to minimize clone. Developers should consider trimming or replacing libraries with high clone rates, especially when only a small subset of features is required. When robust frameworks are indispensable (e.g., Oculus SDK or SteamVR), their integration should be abstracted via dependency injection or interface wrappers to avoid deeply embedded code dependencies and facilitate easier maintenance.
Finally, to address the risks of inter-version cloning, teams should establish cross-version code management workflows. This includes centralizing reusable logic in shared core modules and maintaining traceability of duplicated segments to ensure timely propagation of fixes and enhancements across all project versions.
5.2 Limitation and Future Work
The present study has several limitations that need to be addressed in future research. First, the scope of our dataset presents certain constraints that may affect the generalizability of our findings. (i) Currently, the dataset primarily consists of C# projects, aligning with the current landscape of open-source VR development. Although the exclusion of non-NiCad-supported languages such as JavaScript affects only a small subset of auxiliary tools, extending language support is essential for enhancing generalizability—particularly in multilingual and proprietary VR contexts. (ii) Another key limitation of our study lies in the exclusion of Unreal Engine-based projects, despite their prominence in VR development. This exclusion stems from two primary factors: most Unreal Engine-based VR projects are proprietary and not publicly accessible as open-source code; and such projects are predominantly developed in C++, which is not supported by our clone detection tool, NiCad. These constraints collectively prevent the inclusion of Unreal Engine projects in our dataset and may limit the generalizability of our findings to closed-source VR ecosystems. We suggest that future research explore alternative analysis methods or tools capable of supporting C++ and binary formats, thereby broadening the scope of VR project analysis.
Second, the employment of GPT-4o for file clone detection presents several limitations that warrant discussion. (i) Different versions or configurations of the GPT-4o can sometimes produce varying similarity values, highlighting the need for consistent model configurations and fine-tuning to reduce discrepancies in similarity tasks. (ii) Although our study illustrates the effectiveness of GPT-4o in asset file clone detection for open-source VR software, its scalability and real-time performance in large-scale VR systems have not yet been evaluated, and current LLMs often underperform in such settings. In addition, while fine-tuned LLMs could improve detection precision for serialized assets, the absence of domain-specific datasets currently hinders such efforts. To address these issues, our future work will focus on building dedicated asset file corpora and training customized models. (iii) The application of LLMs in code and file clone detection within VR software introduces ethical concerns, particularly regarding data privacy and inference bias. While our reliance on open-source datasets mitigates these concerns in the present study, future research targeting proprietary VR projects or user-contributed content should adopt more robust privacy-preserving and fairness-aware practices.
Another limitation of this study lies in its absence of VR-specific clone metrics. The adopted metrics are derived from traditional software research and remain overly straightforward. Given the complexity and unique characteristics of VR environments, there exists a pressing need to establish more sophisticated, VR context-aware metrics that correspond to varied analytical objectives, coupled with systematic evaluations encompassing strengths, limitations, verification, and scenario-specific applicability.
Finally, our current evaluation methodology examines source code and serialized asset cloning in isolation. An important research direction would involve developing an integrated assessment framework that simultaneously analyzes both artifact types. This may be achieved through the design of a specialized LLM-based analysis agent capable of cross-artifact pattern recognition, unified clone behavior modeling, and combined metric computation—thereby enabling truly holistic VR software clone analysis.